# Evaluating Intuitive Physics Understanding in Video Diffusion Models via Likelihood Preference

### v1

@@ 4.1 Main results
Moreover, we conduct a human study to assess how well PPE aligns with human judgments of physical consistency in generated videos (see Sec. D.2 for details). For each model, we generated 120 samples using the same prompts that describe the designed physics scenarios. Following the VideoPhy annotation protocol (Bansal et al., 2025), human annotators rated each video on a 1–5 scale, where 1 indicates a severe violation of physics and 5 indicates strong consistency. We then derived a ranking of models from the aggregated human scores and compared it against the PPE-based ranking of LikePhys, measuring their agreement with Kendall’s $\tau$. We further compare PPE with state-of-the-art automatic evaluators, including human-aligned VLM such as VideoPhy (Bansal et al., 2024) and VideoPhy2 (Bansal et al., 2025), as well as Qwen 2.5 VL (Bai et al., 2025) as a general VLM prompted for evaluating physical correctness. We find that PPE achieves a stronger overall correlation with human annotations, demonstrating the competitiveness of our zero-shot, training-free approach. In particular, the resulting correlation of $\tau=0.44$ indicates that models with lower PPE also tend to be judged by humans as more physically consistent.
@@
While we verify the correlation between PPE and physics consistency of generated videos, we examine whether it also correlates with established visual quality metrics from VBench (Huang et al., 2024) reported in Tab. 3. By doing so, we aim to determine whether perceptual metrics already capture physical plausibility or represent an independent dimension of evaluation. As shown in the table, aesthetic quality has no significant correlation (r=0.05r=-0.05) to PPE, indicating our metric targets physical correctness rather than aesthetic appeal. Subject consistency (r=0.01r=-0.01) and background consistency (r=0.01r=-0.01) show very weak correlations, consistent with the fact that the feature similarity-based metric does not fully capture physics correctness of dynamic motion. In contrast, motion smoothness ($r=0.15$) and temporal flickering ($r=0.12$) have moderate positive correlations with PPE. This is expected, as smooth, flicker-free motion can be a partial indicator of physical plausibility. Overall, these results confirm that our method evaluates aspects of physical reasoning that are largely orthogonal to visual quality measures, providing complementary insights into video generative models’ capabilities and failure modes.
@@ 4.2 Analysis on Influencing Factors
Role of Classifier-free Guidance. Classifier-free guidance (CFG) (Ho & Salimans, 2022) is widely used to balance fidelity and diversity in diffusion sampling, but its effect on physics understanding remains unclear. We evaluate models across varying CFG strengths and find that guidance scale has only a marginal impact on physics understanding. This insensitivity suggests that, unlike visual quality metrics, physics plausibility is governed primarily by the learned model distribution, while inference-time calibration of noise prediction plays only a minor role. Consequently, CFG can be tuned for visual quality without compromising the model’s physics understanding.
@@ 4.3 Analysis across physics domains and laws
5 Discussion and Limitation
In LikePhys, by leveraging the violation of expectation paradigm united with the density estimation capacity of video diffusion model, we ensure that any difference in denoising loss arises purely from physical implausibility in controlled video pairs. Since generative models can use different spatial and temporal resolutions and inherit style biases from training data, generation-based metrics often capture appearance noise unrelated to physics. In contrast, we compare density estimates for each controlled pair within the same diffusion model, which removes constant biases and makes the evaluation directly comparable across architectures and model scales. Our approach also has limitations and relies on specific assumptions, that we discuss below.
Training data. A key assumption underlying our approach is the model training data predominantly obey real-world physical laws, since the model learns to approximate the data distribution. We believe this assumption is reasonable, given that widely used dataset such as HD-VILA-100M (Xue et al., 2022), and Panda-70M (Chen et al., 2024) consist mainly of real-world recordings rather than animated or synthetic content. Training on such data should encourage generative models to capture true physical dynamics, forming a prerequisite for video generative models that aim to function as physical world simulators.
Cost of Test Data Curation One limitation of our method is that it relies on controlled physics violations, either within our designed benchmark or in existing intuitive physics datasets. Extending to a broader range of scenarios would require additional data curation, which is more costly compared to text-conditioned generation–based evaluation protocols. On the other hand, our approach offers the unique advantage of precisely controlling valid–invalid pairs to probe specific laws and failure patterns, something that text-conditioned generation cannot guarantee.
Requirement on Accessing Noise Prediction Another limitation is that our method requires access to the noise prediction error of the model, which makes it difficult to evaluate closed-source models (Google DeepMind, 2025; Brooks et al., 2024). However, this requirement is less restrictive in the open source community. More importantly, our proposed method can be used as an indicator for monitoring training progress and selection of checkpoints during model release.
6 Conclusion
We introduced LikePhys, a training-free likelihood-preference evaluation method and benchmark for assessing the intuitive physics understanding of text-to-video diffusion models. Our systematic evaluation shows that while current models still fall short in intuitive physics in complex and chaotic dynamics and exhibit varying performance across different physics domains, clear progress emerges with model scale, training data size, and inference settings, underscoring practical directions toward physics-aware generative modelling.
@@ Appendix C Additional Details of Inference Settings
Our evaluation leverages off‐the‐shelf video diffusion pipelines in a zero‐shot setting. For each model, we adopt the officially recommended spatial and temporal resolutions as shown in Tab. 4 and apply a uniform noise scheduler (Euler discrete) over a fixed number of timesteps ($T=10$) as uniform sampling found optimal in Diffusion Classifier (Li et al., 2023).
@@ D.2 Human Study Experiment Details
For the human study experiment, we generate 10 samples per physics scenario per model using the same prompts as shown in Apx. C with prompt enhancement from GPT5. Human annotators are asked to rate each video on a scale of 1–5, following the annotation protocol of VideoPhy2 (Bansal et al., 2025), where higher scores indicate stronger physical consistency.
For baseline physics evaluator comparison, we leverage VideoPhy1 (Bansal et al., 2024), which rates videos on Physics Consistency (PC) and Semantic Adherence (SA) using a 0–1 scale. Following the official evaluation protocol (Bansal et al., 2024), we obtain per-model scores by filtering out samples with PC $\leq 0.5$ or SA $\leq 0.5$, and then computing the proportion of samples that pass this joint criterion. For VideoPhy2, which rates PC and SA on a 1–5 scale, we follow the official evaluation practice (Bansal et al., 2025) by filtering out samples with Physics Consistency $\leq 4$ or Semantic Adherence $\leq 4$, and then calculating the proportion of valid samples after filtering. For Qwen2.5-VL (Bai et al., 2025), we follow the evaluation protocol of DreamGen (Jang et al., 2025). Each video is rated on a 0–1 scale using the following prompt:
@@ Appendix F Applicability to Other Intuitive Physics Benchmarks
We further demonstrate that LikePhys applies beyond our main setups by evaluating on the IntPhys Development split (Riochet et al., 2018). The Dev split contains three blocks including O1 (object permanence), O2 (shape constancy), and O3 (spatio-temporal continuity). Each constructed as matched pairs of possible against impossible events. We run LikePhys unchanged and report PPE per block and averaged across blocks. As shown in Tab. 6, models exhibit distinct profiles across principles (e.g., some perform better on continuity in O3 than on permanence in O1), mirroring the domain-specific trends we observed on our rigid-body and optics evaluations. This supports the portability of LikePhys and its ability to surface principle-level strengths and weaknesses without task-specific tuning.

### v2

@@ 4.1 Main results
Model ranking
In Tab. 1, we benchmark twelve VDMs and order them by increasing PPE. The results reveal large performance gaps across architectures. Early UNet-based models (ronneberger2015u) such as AnimateDiff and ZeroScope often record error rates above 50%, reflecting limited ability to distinguish physically valid videos from invalid ones. In contrast, more recent DiT-based designs (peebles2023dit), including the top three performers Hunyuan T2V (43.6%), Wan2.1–T2V–14B (43.8%), and CogVideoX1.5–5B (43.8%), achieve substantially lower errors across multiple scenarios (see Sec. 4.2 for further discussion).
Moreover, while there is a clear trend of improvement in recent architectures, only a handful of models significantly achieve PPE lower than the 50% random-guess threshold in more than a few scenarios, potentially due to simplicity biases, where physically implausible interactions appear visually easier to model than valid ones (shah2020pitfalls; geirhos2020shortcut). These results highlight that while modern VDMs start to internalise physical principles, their performance varies substantially across different physics domains (see Sec. 4.3 for further analysis), indicating significant room for progress in physically-accurate training of video generators.
Alignment to Human Preference
Moreover, we conduct a human study to assess how well PPE serves as a proxy and aligns with human judgments of physics plausibility in downstream video generation. Specifically, we perform text-to-video generation with model tested in Sec. 4.1, generating 120 samples per model using the prompts that describe the designed physics scenarios in Sec. 3.3. Following the VideoPhy annotation protocol (bansal2025videophy2), human annotators rate each video on a 1–5 scale, where 1 indicates a severe violation of physics and 5 indicates strong consistency (see Sec. D.2 for details). We then derive a ranking of models from the aggregated human scores and compare it against the PPE-based ranking of LikePhys, measuring their agreement with Kendall’s $\tau$. We note that the VLM-based rankings leverage the dataset of videos generated by the candidate models, whereas the PPE-based ranking is computed solely on our synthetic benchmark and does not use any downstream model-generated videos.
We further compare PPE against state-of-the-art automatic evaluators, including human-aligned VLM-based metrics such as VideoPhy (bansal2024videophy) and VideoPhy2 (bansal2025videophy2), as well as Qwen 2.5 VL 7B-Instruct (bai2025qwen2; jang2025dreamgen) as a general VLM prompted to evaluate physical correctness. We find that PPE achieves a stronger overall correlation with human annotations, demonstrating the competitiveness of our zero-shot, training-free approach. In particular, the resulting correlation of $\tau=0.44$ indicates that models with lower PPE also tend to be judged by humans as more physically consistent.
@@
While we verify the correlation between PPE and physics consistency of generated videos, we examine whether it also correlates with established visual quality metrics from VBench (huang2024vbench) reported in Tab. 3. By doing so, we aim to determine whether perceptual metrics already capture physical plausibility or represent an independent dimension of evaluation. As shown in the table, aesthetic quality has no significant correlation (r=0.05r=-0.05) to PPE, indicating our metric targets physical correctness rather than aesthetic appeal. Subject consistency (r=0.01r=-0.01) and background consistency (r=0.01r=-0.01) show very weak correlations, consistent with the fact that the feature similarity-based metric does not fully capture physics correctness of dynamic motion. In contrast, motion smoothness ($r=0.15$) and temporal flickering ($r=0.12$) have moderate positive correlations with PPE. This is expected, as smooth, flicker-free motion can be a partial indicator of physical plausibility. Overall, these results confirm that our method evaluates aspects of physical reasoning that are largely orthogonal to visual quality measures, providing complementary insights into video generative models’ capabilities and failure modes.
@@ 4.2 Analysis on Influencing Factors
Role of Classifier-free Guidance. Classifier-free guidance (CFG) (ho2022classifier) is widely used to balance fidelity and diversity in diffusion sampling, but its effect on physics understanding remains unclear. We evaluate models across varying CFG strengths and find that guidance scale has only a marginal impact on physics understanding. This insensitivity suggests that, unlike visual quality metrics, physics plausibility is governed primarily by the learned model distribution, while inference-time calibration of noise prediction plays only a minor role. Consequently, CFG can be tuned for visual quality without compromising the model’s physics understanding.
@@ 4.3 Analysis Across Physics Domains and Laws
5 Discussion and Conclusion
In this paper, we introduce LikePhys, a method for evaluating physical understanding in video diffusion models. By leveraging the violation-of-expectation paradigm together with the density estimation capability of video diffusion models, we ensure that any difference in denoising loss within a controlled video pair arises purely from physical implausibility. While our studies show the effectiveness of LikePhys in evaluating physics understanding, our approach also has limitations. We further discuss them and open new doors for research.
Empirical Assessment of Physics Understanding. Our method assesses whether the distribution learned by a video generative model is close to a physics-plausible distribution by comparing the likelihoods the model assigns to physically valid and invalid video pairs. Several factors can influence this physics understanding capacity, such as the training data distribution, suboptimal optimisation, and the training objective. However, we do not make any prior assumption about the cause of this capacity. Instead, LikePhys provides an empirical assessment of the resulting physics understanding.
Cost of data curation. One limitation of our method is that it relies on controlled physics violations, which is impossible to obtain in real-world. Extending to a broader range of scenarios would require additional data curation in simulator or video editing. While this is more costly compared to text-conditioned generation–based evaluation protocols, our approach offers the unique advantage of precisely controlling valid–invalid pairs to probe specific laws and failure patterns.
Accessing Noise Prediction Our method requires access to the noise prediction error of the model, which makes it difficult to evaluate closed-source models (deepmind_veo3; sora). However, this requirement is less restrictive in the open source community. We believe that, during the development of closed-source models, our proposed method can be used as an indicator for monitoring training progress and selecting checkpoints during model release, which may require ranking hundreds of models.
@@ Appendix C Additional Details of Inference Settings
Our evaluation leverages off‐the‐shelf video diffusion pipelines in a zero‐shot setting. For each model, we adopt the officially recommended spatial and temporal resolutions as shown in Tab. 4 and apply a uniform noise scheduler (Euler discrete) over a fixed number of timesteps ($T=10$) as uniform sampling found optimal in Diffusion Classifier (li2023diffusion).
@@ D.2 Human Study Experiment Details
For the human study experiment, we generate 10 samples per physics scenario per model using the same prompts as shown in Apx. C with prompt enhancement from GPT5. Human annotators are asked to rate each video on a scale of 1–5, following the annotation protocol of VideoPhy2 (bansal2025videophy2), where higher scores indicate stronger physical consistency.
For baseline physics evaluator comparison, we leverage VideoPhy1 (bansal2024videophy), which rates videos on Physics Consistency (PC) and Semantic Adherence (SA) using a 0–1 scale. Following the official evaluation protocol (bansal2024videophy), we obtain per-model scores by filtering out samples with PC $\leq 0.5$ or SA $\leq 0.5$, and then computing the proportion of samples that pass this joint criterion. For VideoPhy2, which rates PC and SA on a 1–5 scale, we follow the official evaluation practice (bansal2025videophy2) by filtering out samples with Physics Consistency $\leq 4$ or Semantic Adherence $\leq 4$, and then calculating the proportion of valid samples after filtering. For Qwen2.5-VL (bai2025qwen2), we follow the evaluation protocol of DreamGen (jang2025dreamgen). Each video is rated on a 0–1 scale using the following prompt:
@@ Appendix F Applicability to Other Intuitive Physics Benchmarks
We further demonstrate that LikePhys applies beyond our main setups by evaluating on the IntPhys Development split (riochet2018intphys). The Dev split contains three blocks including O1 (object permanence), O2 (shape constancy), and O3 (spatio-temporal continuity). Each constructed as matched pairs of possible against impossible events. We run LikePhys unchanged and report PPE per block and averaged across blocks. As shown in Tab. 6, models exhibit distinct profiles across principles (e.g., some perform better on continuity in O3 than on permanence in O1), mirroring the domain-specific trends we observed on our rigid-body and optics evaluations. This supports the portability of LikePhys and its ability to surface principle-level strengths and weaknesses without task-specific tuning.